# Candidate Monotonicity and Proportionality for Lotteries and Non-Resolute Rules

### v1

@@ 1 Proportionality and Monotonicity
Proportional multiwinner voting rules are among the most well studied in computational social choice. Among them, the Single Transferable Vote (STV) is used in several countries, e.g., Australia, Malta, or Scotland, to elect governing bodies. While STV can be considered proportional (Dummett, 1984; Tideman, 1995) one of its flaws is its failure of candidate monotonicity: strengthening a winning candidate can make this candidate stop winning (Woodall, 1997; McCune and Graham-Squire, 2024). Interestingly enough, whether this flaw is a necessity is an open question. Namely, it is unknown whether there is any voting rule satisfying the proportionality axiom of proportionality for solid coalitions (PSC) together with candidate monotonicity (Woodall, 1997; Aziz and Lee, 2020) with most known proportional voting rules failing candidate monotonicity already for very small instances sizes (Aziz and Lee, 2020). Despite Woodall already writing that “a major unsolved problem is whether there exist rules that retain the important political features of STV and are also more monotonic” (Woodall, 1997) no such rule has been found or proven to be nonexistent.
Our Contribution.
@@
In this paper, we tackle this open question from two slightly weaker angles. Firstly, we study non-resolute voting rules: rules that do not just output one committee, but a whole set of committees. For this setting, we show that the rule outputting all committees satisfying PSC is indeed candidate monotone. Secondly, we study probabilistic voting rules, i.e., rules selecting a lottery over committees. For this, we show that a variant of Phragmén’s Ordered Rule (Janson, 2016) can be adapted into a candidate monotone lottery over committees satisfying PSC.
Related Work.
@@
Our work is situated in the area of proportional multiwinner voting with ranked preferences, in particular in the axiomatic study of multiwinner voting rules. This study dates back to the work of eminent philosopher Dummett (1984) and his development of PSC. Aziz and Lee (2020) were the first to give slightly improved results for monotonicity and proportionality, showing that their Expanding Approvals Rule (EAR) satisfies a weaker form of monotonicity together with a generalization of PSC in the setting of weak ranked preferences in which ties are allowed. The same generalization of PSC together with a form of monotonicity they call “indifference monotonicity” is also satisfied by the generalization of STV by Delemazure and Peters (2024). McCune and Graham-Squire (2024) experimentally studied monotonicity violations in a real-world dataset consisting of Scottish local elections and found that out of 1079 elections 62 possessed some kind of “monotonicity anomaly”. In particular, in $23$ elections STV failed candidate monotonicity as studied by us.
@@
PSC was further studied by Brill and Peters (2023) who criticized it for being a relatively weak axiom and introduced stronger proportionality axioms still satisfied by EAR. Further, Aziz and Lee (2021) introduced PSC generalizations for the context of participatory budgeting (Peters et al., 2021) and in Aziz and Lee (2022) provided a characterization of PSC in the standard ordinal preference setting. Finally, recently Aziz et al. (2024a) studied committee monotonicity, i.e., the selection of a ranking of candidates, together with PSC and showed that a rule satisfying both PSC and committee monotonicity always exists.
@@
In the setting of approval-based multiwinner voting, monotonicity is significantly easier to obtain (Sánchez-Fernández and Fisteus, 2019) with most rules being monotone in the sense that voters additionally approving an already winning candidate cannot hurt that candidate. However, it can for instance happen that adding voters only approving a single candidate can make this candidate not be chosen anymore (see Boehmer et al. (2024) for examples of this in the context of participatory budgeting).
@@
Finally, our results are also situated in the area of probabilistic social choice. Here, we are closely related to recent work focusing on best of both world guarantees, i.e., the design of lotteries satisfying desirable properties both ex-ante and ex-post. In particular, recent work focusing on the design of randomized apportionment schemes (Gölz et al., 2024; Correa et al., 2024) and on achieving ex-post and ex-ante proportionality guarantees in approval-based multiwinner voting and participatory budgeting (Aziz et al., 2023, 2024b; Suzuki and Vollen, 2024) are close to our section on probabilistic voting rules.
@@ 2 Preliminaries
Notation.
For $t\in\mathbb{N}$ let $[t]\coloneqq\{1,\dots,t\}$. We are given a set $N=[n]$ of voters and a set $C=\{c_{1},\dots,c_{m}\}$ of candidates. Each voter $i\in N$ possesses a strict linear ranking $\succ_{i}\subseteq C\times C$ over all candidates. Further, we are given a committee size $k\leq m$. Together, the voters, candidates, preferences, and committee size form an instance $\mathcal{I}=(N,C(\succ_{i})_{i\in N},k)$. A committee is a subset $W\subseteq C$ of size $\lvert W\rvert=k$. Let $C_{k}=\{W\subseteq C\colon\lvert W\rvert=k\}$ be the set of all committees. A rule $r$ is a function mapping an instance $\mathcal{I}$ to a non-empty set of committees $r(\mathcal{I})\subseteq C_{k}$. If $\lvert r(\mathcal{I})\rvert=1$ for all instances $\mathcal{I}$ the rule is said to be resolute.
For probabilistic rules, we mostly follow the notation of Aziz et al. (2023). We say that a fractional committee is an assignment of probabilities $(p_{1},\dots,p_{m})\in[0,1]^{m}$ for each candidate with $\sum_{i\in m}p_{i}=k$ with a fractional committee rule mapping instances to fractional committees. We let $\Delta(C_{k})$ be the set of all probability distributions over $C_{k}$. A probabilistic rule $r$ then maps an instance $\mathcal{I}$ to a probability distribution $r(\mathcal{I})\in\Delta(C_{k})$. Each probabilistic rule also has a natural non-resolute counterpart $\operatorname{supp}(r(\mathcal{I}))$, i.e., the set of committees being selected with non-zero probability (or the support of $r(\mathcal{I})$). Further, each probabilistic rule $r$ implements a fractional committee voting rule, with the probability of each candidate being the probability of this candidate being on the committee selected by $r$. On the other hand, there can be several probabilistic voting rules implementing a fractional committee voting rule.
Proportionality.
Following the seminal work of Dummett (1984) we define the axiom of proportionality for solid coalitions (PSC). For this, we first need to define the eponymous solid coalitions.
@@
$C^{\prime}\succ_{i}C\setminus C^{\prime}\text{ for all }i\in N^{\prime}.$
@@
Definition 2 (Proportionality for Solid Coalitions (Dummett, 1984)).
@@
    $\lvert W\cap C^{\prime}\rvert<\min(\lvert C^{\prime}\rvert,\ell)$.
@@
There are several rules satisfying PSC, among them the famous single-transferable vote (STV) (Tideman, 1995) and the expanding approvals rule (EAR) (Aziz and Lee, 2020).
Note that due to our definition of rules the set of all committees satisfying PSC is also a (non-resolute) rule.
@@
This instance contains one non-trivial solid coalition large enough to be able to demand a candidate, namely the first and second voter over $\{c_{1},c_{2}\}$ (and technically also over $\{c_{1},c_{2},c_{3}\}$). Thus, any committee containing at least one one $c_{1}$ or $c_{2}$ satisfies PSC. Now if the third voter would swap candidate $c_{4}$ up the third and fourth voter would form a solid coalition over $\{c_{4}\}$. Therefore, in this new instance the only two PSC committees are $\{c_{1},c_{4}\}$ and $\{c_{2},c_{4}\}$
Monotonicity.
For our definition of (candidate) monotonicity for non-resolute voting rules, we follow Elkind et al. (2017).
Definition 3 (Candidate Monotonicity (Elkind et al., 2017)).
@@
That is, if the candidate $c$ is in some winning committee and is “made stronger” $c$ is still in some winning committee afterward. It is known that neither STV nor EAR are candidate monotone (Aziz and Lee, 2020). However, Phragmén’s ordered rule is indeed candidate monotone (Janson, 2016).
@@ 3 Non-Resoluteness and Monotonicity
We start off our results by showing that the rule that selects all PSC committees is candidate monotone. For this we crucially make use of a characterization of PSC by Aziz and Lee (2022) using their minimal demand rules. These works as follows: We start off with an empty committee $W=\emptyset$. For each rank $r=1,\dots,m$ we check if there is a group of voters forming a solid coalition over a set of candidates of size $r$. If, additionally, this solid coalition witnesses a PSC violation, we add a candidate from $C^{\prime}$ to $W$. We repeat this until there is no further PSC violation, afterwards we increment $r$.
A committee now satisfies PSC if and only if it can be obtained via this process.
Theorem 2 (Characterization of PSC (Aziz and Lee, 2022)).
@@
Let $W$ be a committee satisfying PSC and $c\in W$. Assume that there is voter $i$ swapping $c$ up in their preference list. As $W$ satisfies PSC in the original preferences, it can be obtained via a minimal demand rule. If $W$ still satisfies PSC, even after the preference change, we are already done. Otherwise, there must exist a solid coalition $N^{\prime}$ and $C^{\prime}$ witnessing a PSC violation. We now follow the characterization of Aziz and Lee (2020) and iteratively check for each rank if there is a solid coalition witnessing a PSC violation. If there is one and the solid coalition does not include $i$ and $c$ this solid coalition already exists in the original instance. Thus, this solid coalition must have appeared in the original minimal demand rule, and we can include the same candidate as was included in $W$. If the solid coalition includes $i$ and $c$ however, we can include $c$ as it is part of the candidates witnessing the violation. As $W$ does not satisfy PSC, such a violation must appear eventually. The rest of the committee can then be filled up with an arbitrary execution of the minimal demand rule, therefore leading to a committee satisfying PSC and including $c$. ∎
@@ 4 A Probabilistic Version of Phragmén’s Ordered Rule
While the previous result gives us some hope for obtaining a proportional and candidate monotone rule, it is also quite weak. There can be a lot of committees satisfying PSC (Bardal et al., 2024) and our monotonicity notion only guarantees us that some committee includes the strengthened candidate. To remedy this, we study probabilistic voting rules. In particular, we use a variant of the Phragmén’s Ordered Rule (Janson, 2016) and turn it into a lottery over committees satisfying PSC.
Phragmén’s Ordered Rule.
Phragmén’s Ordered Rule was originally introduced by a Swedish royal commission (including Phragmén himself) and was used for the intra party seat distribution in the Swedish parliament, for more historic context see the work by Janson (2016).
Phragmén’s Ordered Rule works as a continuous process. Each voter starts to “eat” their top choice candidate at the same speed. Once a candidate is fully eaten, this candidate is elected onto the committee and the voters move on to their top choice “uneaten” candidate. This is done until $k$ candidates are elected. [^2] This is very similar to the famous probabilistic serial method of Bogomolnaia and Moulin (2001) and the veto by consumption rule of Ianovski and Kondratev (2021). For a connection between these three methods and the problem of designing a voting rule with low metric distortion see the note by Peters (2023).
@@
    • The first two voters move to their next choices $c_{4}$ ad $c_{3}$.
@@
    • $c_{3}$ is now fully eaten by the second and fourth voter, $c_{3}$ is $75\%$ eaten and $c_{4}$ is $25\%$ eaten.
@@
In the instance in Example 3, $c_{6}$ would get a probability of $1$ while $c_{1}$ to $c_{5}$ each get a probability of $\frac{2}{5}$.
While PFR gives us probabilities, it does not actually give us a distribution over committees yet. For this, we turn to the dependent rounding scheme of Gandhi et al. (2006).
Theorem 4 ((Gandhi et al., 2006)).
@@
    $\sum_{c\in C^{\prime}}p_{c}<\min(\lvert C^{\prime}\rvert,\ell)$.
Now we turn to the main result of this section and show that PFR is indeed candidate monotone and satisfies this ex-ante PSC notion. Further, we use the dependent rounding scheme of Gandhi et al. (2006) to show that the probabilities given by PFR can be decomposed into a probabilistic rule only randomizing over committees satisfying PSC.
@@
The monotonicity of PFR follows from the proof of Janson (2016, Theorem 14.3), as he shows that at any time, after a candidate gets swapped up, this candidate is always “at least as much eaten” as before swapping this candidate up. In particular this holds for time step $k$ and thus the monotonicity follows.
@@
For the ex-ante PSC guarantee, consider any group of size $\ell\frac{n}{k}$ forming a solid coalition over $C^{\prime}$. After $k$ time steps, the group has eaten exactly $\ell$ parts of candidates. As $C^{\prime}$ forms a prefix of the preferences of these voters, it must hold that either $C^{\prime}$ is eaten in total, thus giving a probability mass of $\lvert C^{\prime}\rvert$ or that at least $\ell$ probability mass is spent on candidates in $C^{\prime}$ as voters in $N^{\prime}$ would first start eating candidates outside of $C^{\prime}$ once everything in $C^{\prime}$ is eaten.
@@
For the ex-post guarantee, we make use of the dependent rounding scheme of Theorem 4 while following the general structure of the minimal demand rules of Aziz and Lee (2022).
@@
For a candidate $c\in C$ and voter $i\in N$ let $p_{c}(i)\in[0,1]$ be the amount of $c$ eaten by $i$ during the execution of PFR. Further, let $p_{c}$ be the probability of candidate $c$ being selected by PFR.
@@
We start off with the bipartite graph $G=(\emptyset\cup C,E)$. As a first step, for each candidate $c\in C$ with $p_{c}=1$ we add a voter $v_{c}$ to $G$ with an edge to $c$ of weight $1$ and set $p_{c}(i)=0$ for all $i\in V$. Now, we go from rank $r=1$ to $m$. If there is a group $N^{\prime}\subseteq N$ of voters forming a solid coalition with $C^{\prime}$ of size $\lvert C^{\prime}\rvert=r$, let $\ell=\left\lfloor\sum_{i\in N^{\prime}}\sum_{c\in C^{\prime}}p_{c}(i)\right\rfloor$. If $\ell=0$ we continue. Otherwise, let $p^{\prime}_{c}(i)$ for $c\in C^{\prime}$ and $i\in N^{\prime}$ such that $p^{\prime}_{c}(i)\leq p_{c}(i)$ and $\sum_{c\in C^{\prime},i\in N^{\prime}}p^{\prime}_{c}(i)=\ell$. We add a group voter $i^{\prime}$ to the left side of the bipartite graph and connect it to each candidate in $c\in C^{\prime}$ with probability $\sum_{i\in N^{\prime}}p^{\prime}_{c}(i)$. Further, we deduct $p_{c}(i)$ by $p^{\prime}_{c}(i)$. If there is no such group, we increase $r$.
The dependent rounding scheme now selects committees according to $G$.

### v2

@@ Abstract
keywords:
Computational Social Choice, Multiwinner Voting, Monotonicity
@@ 1 Proportionality and Monotonicity
Proportional multiwinner voting rules are among the most well studied in computational social choice. Among them, the Single Transferable Vote (STV) is used in several countries, e.g., Australia, Malta, or Scotland, to elect governing bodies. One of the main reasons STV can be considered proportional is its satisfaction of the proportionality for solid coalitions (PSC) axiom [Dumm84a, Tide95a]. In essence, for ordinal preferences lists PSC requires that every group of voters listing the same block of candidates (in potentially different permutations) at the start of their preference list, should get proportionally many candidates from this block. For instance, if $5$ candidates were to be elected and $40\%$ of the voters have the same three candidates in their top three, then two out of these three candidates need to be chosen in the end. [^1] This is slightly simplifying things, the commonly used variants of STV usually work with a “quota” that is slightly lower, the so-called Droop quota $\lceil\frac{n}{k+1}\rceil$ with $n$ being the number of voters and $k$ the number of candidates to be elected. For ease of exposition, we focus on the Hare quota $\frac{n}{k}$ instead. The results also hold for the Droop quota. One of the major flaws of STV is its failure of candidate monotonicity: strengthening a winning candidate can make this candidate stop winning [Wood97a, McGr23a]. Interestingly enough, whether this flaw is a necessity is an open question. Namely, it is unknown whether there is any voting rule satisfying PSC together with candidate monotonicity [Wood97a, AzLe20a] with most known proportional voting rules failing candidate monotonicity already for very small instance sizes [AzLe20a]. Despite Wood97a already writing that “a major unsolved problem is whether there exist rules that retain the important political features of STV and are also more monotonic” [Wood97a, page 1] no such rule has been found or proven to be nonexistent.
Our Contribution
@@
In this paper, we tackle this open question from two slightly weaker angles. Firstly, we study non-resolute voting rules: rules that do not just output one committee, but a whole set of committees. For this setting, we show that the rule outputting all committees satisfying PSC is indeed candidate monotone. Secondly, we study probabilistic voting rules, i.e., rules selecting a lottery over committees. For this, we show that a variant of Phragmén’s Ordered Rule [Jans16a] can be adapted into a candidate monotone lottery over committees satisfying PSC.
Related Work
@@
Our work is situated in the area of proportional multiwinner voting with ranked preferences, in particular in the axiomatic study of multiwinner voting rules. This study dates back to the work of eminent philosopher Dumm84a and his development of PSC. AzLe20a were the first to give slightly improved results for monotonicity and proportionality, showing that their Expanding Approvals Rule (EAR) satisfies a weaker form of monotonicity together with a generalization of PSC in the setting of weak ranked preferences in which ties are allowed. The same generalization of PSC together with a form of monotonicity they call “indifference monotonicity” is also satisfied by the generalization of STV by DePe24a. McGr23a experimentally studied monotonicity violations in a real-world dataset consisting of Scottish local elections and found that out of 1079 elections 62 possessed some kind of “monotonicity anomaly”. In particular, in $23$ elections STV failed the version of candidate monotonicity studied by us.
@@
PSC was further investigated by BrPe23a who criticized it for being a relatively weak axiom and introduced stronger proportionality axioms still satisfied by EAR. Further, AzLe21a introduced PSC generalizations for the context of participatory budgeting [PPS21a] and in AzLe22a provided a characterization of PSC in the standard ordinal preference setting. Finally, recently ALR24a studied committee monotonicity, i.e., the selection of a ranking of candidates, together with PSC and showed that a rule satisfying both PSC and committee monotonicity exists.
@@
In the setting of approval-based multiwinner voting, monotonicity is significantly easier to obtain [SaFi17a] with most rules being monotone in the sense that voters additionally approving an already winning candidate cannot hurt that candidate. [^2] In fact, all voting rules studied in the recent book of LaSk22a on approval-based multiwinner voting satisfy candidate monotonicity [LaSk22a, Table 3.1] However, it can for instance happen that adding voters only approving a single candidate can make this candidate not be chosen anymore (see BFJ+24a for examples of this in the context of participatory budgeting).
@@
Finally, our results are also situated in the area of probabilistic social choice. Here, we are closely related to recent works focusing on best of both world guarantees, i.e., the design of lotteries satisfying desirable properties both ex-ante and ex-post. In particular, recent work focusing on the design of randomized apportionment schemes [GPP24a, JGU+24a] and on achieving ex-post and ex-ante proportionality guarantees in approval-based multiwinner voting and participatory budgeting [ALS+23a, ALS+24a, SuVo24a] are close to our section on probabilistic voting rules.
@@ 2 Preliminaries
Notation
For $t\in\mathbb{N}$ let $t]\coloneqq\{1,\dots,t\}$. We are given a set $N=[n]$ of voters and a set $C=\{c_{1},\dots,c_{m}\}$ of candidates. Each voter $i\in N$ possesses a strict linear ranking $\succ_{i}\subseteq C\times C$ over all candidates. Further, we are given a committee size $k\leq m$. Together, the voters, candidates, preferences, and committee size form an instance $\mathcal{I}=(N,C,(\succ_{i})_{i\in N},k)$. A committee is a subset $W\subseteq C$ of size $\lvert W\rvert=k$. Let $C_{k}=\{W\subseteq C\colon\lvert W\rvert=k\}$ be the set of all committees. A rule $r$ is a function mapping an instance $\mathcal{I}$ to a non-empty set of committees $r(\mathcal{I})\subseteq C_{k}$. If $\lvert r(\mathcal{I})\rvert=1$ for all instances $\mathcal{I}$ the rule is said to be resolute.
For probabilistic rules, we mostly follow the notation of ALS+23a. We say that a fractional committee is an assignment of probabilities $(p_{1},\dots,p_{m})\in[0,1]^{m}$} for each candidate with $\sum_{i\in[m]}p_{i}=k$ with a fractional committee rule mapping instances to fractional committees. We let $\Delta(C_{k})$ be the set of all probability distributions over $C_{k}$. A probabilistic rule $r$ then maps an instance $\mathcal{I}$ to a probability distribution $r(\mathcal{I})\in\Delta(C_{k})$. Each probabilistic rule also has a natural non-resolute counterpart $\operatorname{supp}(r(\mathcal{I}))$, i.e., the set of committees being selected with non-zero probability (or the support of $r(\mathcal{I})$). Further, each probabilistic rule $r$ implements a fractional committee voting rule, with the probability of each candidate being the probability of this candidate being on the committee selected by $r$. On the other hand, there can be several probabilistic voting rules implementing a fractional committee voting rule.
Proportionality
Following the seminal work of Dumm84a we define the axiom of proportionality for solid coalitions (PSC). For this, we first need to define the eponymous solid coalitions.
@@
[…]
@@
Definition 2 (Proportionality for Solid Coalitions [Dumm84a]).
@@
[…]
@@
There are several rules satisfying PSC, among them the famous Single Transferable Vote (STV) [Tide95a] and the Expanding Approvals Rule (EAR) [AzLe20a]. Note that due to our definition of rules the set of all committees satisfying PSC is also a (non-resolute) rule.
@@
This instance contains one non-trivial solid coalition large enough to be able to demand a candidate, namely the first and second voter over $\{c_{1},c_{2}\}$ (and technically also over $\{c_{1},c_{2},c_{3}\}$). Thus, any committee containing at least one of $c_{1}$ or $c_{2}$ satisfies PSC. Now if the third voter would swap candidate $c_{4}$ up, the third and fourth voter would form a solid coalition over $\{c_{4}\}$. Therefore, in this new instance the only two PSC committees are $\{c_{1},c_{4}\}$ and $\{c_{2},c_{4}\}$
Monotonicity
For our definition of (candidate) monotonicity for non-resolute voting rules, we follow EFSS17a.
Definition 3 (Candidate Monotonicity [EFSS17a]).
@@
That is, if the candidate $c$ is in some winning committee and is “made stronger” $c$ is still in some winning committee afterward. It is known that neither STV nor EAR are candidate monotone [AzLe20a]. However, Phragmén’s ordered rule is indeed candidate monotone [Jans16a].
@@ 3 Non-Resoluteness and Monotonicity
We start off our results by showing that the rule that selects all PSC committees is candidate monotone. For this we crucially make use of a characterization of PSC by AzLe22a using their minimal demand rules. These works as follows: We start off with an empty committee $W=\emptyset$. For each rank $r=1,\dots,m$ we check if there is a group of voters forming a solid coalition over a set $C^{\prime}$ of candidates of size $r$. If, additionally, this solid coalition witnesses a PSC violation, we add a candidate from $C^{\prime}\setminus W$ to $W$. We repeat this until there is no further PSC violation, afterwards we increment $r$. A committee now satisfies PSC if and only if it can be obtained via this process.
Theorem 2 (Characterization of PSC [AzLe22a]).
@@
Let $\mathcal{I}$ be an instance, $W$ a committee satisfying PSC in $\mathcal{I}$, and $c\in W$. Assume that there is a voter swapping $c$ up in their preference list and denote this instance by $\mathcal{I}_{c}$. We now construct a new committee $W_{c}$ containing $c$ by running a minimal demand rule on $\mathcal{I}_{c}$. We start off by initializing $W_{c}=\emptyset$ and iteratively check for each rank $r=1,\dots,m$ if there exists a solid coalition consisting of a set $N^{\prime}\subseteq N$ of voters and $C^{\prime}\subseteq C$ of candidates witnessing a PSC violation in $\mathcal{I}_{c}$. If $c$ is already a part of $W_{c}$ we can include an arbitrary candidate from $C^{\prime}\setminus W_{c}$ to $W_{c}$. Otherwise, we distinguish two cases.
Case 1: $c\notin C^{\prime}$. If this is the case, as $c$ was swapped up, the set $N^{\prime}$ together with $C^{\prime}$ already formed a solid coalition in the original instance $\mathcal{I}$. Thus, as $W$ satisfies PSC there must be more candidates in $W\cap C^{\prime}$ than in $W_{c}\cap C^{\prime}$. Hence, there exists a candidate $c^{\prime}\in(W\cap C^{\prime})\setminus W_{c}$ and we can include this candidate in the committee $W_{c}$.
Case 2: $c\in C^{\prime}$. Then we include $c$ in the committee $W_{c}$.
This constitutes a valid execution of a minimal demand rule. Thus, the committee $W_{c}$ satisfies PSC. If we ever entered the second case, we know that $c\in W_{c}$ has to hold. Otherwise, we always entered the first case. However, then the $k$ candidates selected by $W_{c}$ are precisely the $k$ candidates in $W$. As $c\in W$ this implies $c\in W_{c}$. ∎
@@ 4 A Probabilistic Version of Phragmén’s Ordered Rule
While the previous result gives us some hope for obtaining a proportional and candidate monotone rule, the result is also quite weak. There can be a lot of committees satisfying PSC [BBMP25a] and our monotonicity notion only guarantees us that some committee includes the strengthened candidate. To remedy this, we study probabilistic voting rules. In particular, we use a variant of Phragmén’s Ordered Rule [Jans16a] and turn it into a lottery over committees satisfying PSC.
Phragmén’s Ordered Rule
Phragmén’s Ordered Rule was originally introduced by a Swedish royal commission (including Phragmén himself) and was used for the intra party seat distribution in the Swedish parliament. For more historic context see the work by Jans16a.
Phragmén’s Ordered Rule works as a continuous process. Each voter starts to “eat” their top choice candidate at the same speed. Once a candidate is fully eaten, this candidate is elected onto the committee and the voters move on to their top choice “uneaten” candidate. This is done until $k$ candidates are elected. [^3] This is very similar to the famous probabilistic serial method of BoMo01a and the veto by consumption rule of IaKo21a. For a connection between these three methods and the problem of designing a voting rule with low metric distortion see the note by Pete23a.
@@
    • The first two voters move to their next choices $c_{4}$ and $c_{3}$.
@@
    • $c_{3}$ is now fully eaten by the second and fourth voter, $c_{2}$ is $75\%$ eaten and $c_{4}$ is $25\%$ eaten.
@@
In the instance in Example˜3, $c_{6}$ would get a probability of $1$ while $c_{1}$ to $c_{5}$ each get a probability of $\frac{2}{5}$.
While PFR gives us probabilities, it does not actually give us a distribution over committees yet. For this, we turn to the dependent rounding scheme of GKPS06a.
Theorem 4 ([GKPS06a]).
@@
[…]
To better contexualize ex-ante PSC, we note that a (deterministic) committee satisfying ex-post PSC does not necessarily need to satisfy ex-ante PSC. [^4] As a simple example consider an instance with two voters, one ranking $c_{1}\succ c_{2}$ and the other ranking $c_{2}\succ c_{1}$. Here, for $k=1$, either $\{c_{1}\}$ or $\{c_{2}\}$ satisfies ex-post PSC. However, to satisfy ex-ante PSC, both $c_{1}$ and $c_{2}$ need to be selected with probability $\frac{1}{2}$.
With this definition we now come to the main result of this section and show that PFR is indeed candidate monotone and satisfies this ex-ante PSC notion. Further, we use the dependent rounding scheme of GKPS06a to show that the probabilities given by PFR can be decomposed into a probabilistic rule only randomizing over committees satisfying PSC.
@@
The monotonicity of PFR follows from the proof of Jans16a, as he shows that at any time, after a candidate gets swapped up, this candidate is always “at least as much eaten” as before swapping this candidate up. In particular this holds for time step $k$ and thus the monotonicity follows.
@@
For the ex-ante PSC guarantee, consider any group $N^{\prime}\subseteq N$ of voters of size at least $\ell\frac{n}{k}$ forming a solid coalition over a set $C^{\prime}\subseteq C$ of candidates. After a total weight of $k$ has been eaten, the weight of the candidates eaten by the group $N^{\prime}$ is exactly $\ell$. As $C^{\prime}$ forms a prefix of the preferences of these voters, it must hold that either $C^{\prime}$ is eaten in total or that at least $\ell$ probability mass is spent on candidates in $C^{\prime}$ as voters in $N^{\prime}$ would first start eating candidates outside of $C^{\prime}$ once everything in $C^{\prime}$ is eaten. Thus, the outcome satisfies ex-ante PSC.
@@
For the ex-post guarantee, we make use of the dependent rounding scheme of Theorem˜4 while following the general structure of the minimal demand rules of AzLe22a.
@@
Let $p$ be the fractional committee selected by PFR. For a candidate $c\in C$ and voter $i\in N$ let $p_{c}(i)\in[0,1]$ be the amount of $c$ eaten by $i$ during the execution of PFR.
@@
We start off with the bipartite graph $G=(\emptyset\cup C,E)$ which we will later use to round the outcome. As a first step, for each candidate $c\in C$ with $p_{c}=1$ we add a voter $v_{c}$ to $G$ with an edge to $c$ of weight $1$ and set $p_{c}(i)=0$ for all $i\in N$. Note that if we round $G$ according to the dependent rounding scheme, this forces the edge from $v_{c}$ to $c$ to always be selected. Now, we go from rank $r=1$ to $m$. If there is a group $N^{\prime}\subseteq N$ of voters forming a solid coalition with $C^{\prime}$ of size $\lvert C^{\prime}\rvert=r$, let $\ell=\left\lfloor\sum_{i\in N^{\prime}}\sum_{c\in C^{\prime}}p_{c}(i)\right\rfloor$ be the rounded “budget” left by this group of voters. If $\ell=0$ we continue. Otherwise, for each $c\in C^{\prime}$ and $i\in N^{\prime}$ we select a value $p^{\prime}_{c}(i)\in[0,1]$ such that $p^{\prime}_{c}(i)\leq p_{c}(i)$ and $\sum_{c\in C^{\prime},i\in N^{\prime}}p^{\prime}_{c}(i)=\ell$. We add a group voter $i^{\prime}$ to the left side of the bipartite graph and connect it to each candidate in $c\in C^{\prime}$ with an edge of weight $\sum_{i\in N^{\prime}}p^{\prime}_{c}(i)$. Further, we deduct each $p_{c}(i)$ by $p^{\prime}_{c}(i)$. If there is no group left, we increase $r$.
The dependent rounding scheme now selects committees according to the bipartite graph $G$. That is, we round the edges and select every candidate adjacent to an edge in the rounded outcome. As the weight adjacent to each group voter is integral, and as the total sum of weights is $k$ this guarantees that exactly $k$ candidates are selected. As the weight adjacent to each candidate is at most $1$, we also do not select any candidate twice.